two angles are complimentry. The sum of the measure of the first angle and half the second angle is 69degrees. Find the measures of the angles
Answer by dfrazzetto(283)

2 angles are complementary is their sum is 90 degrees, so:
a + b = 90
a + b/2 = 69 system of equations, multiply bottom by -1
a + b = 90
-a -b/2 = -69 add
_________________
b - b/2 = 21
b/2 = 21
b = 21(2) = 42
a + 42 = 90
a = 90 - 42 = 48
Check:
a + b/2 = 69
48 + 42/2 = 69
48 + 21 = 69
69=69 √